OpenStudy (anonymous):

hi everybody i have a question. the cancer death rate 100,000 people in the United States can be aproximated by the equation y=-2.2x+216, where x is the number of year since 1990. use this equation to determine in whixh year the approximate cancer death rate will be 150?

OpenStudy (anonymous):

no ones knows ?

OpenStudy (anonymous):

y = -2.2x + 216 2.2x = 216 - y x = (216-y) /2.2 x = (216-150)/2.2 x = 30 so in 2020